5- A national test is held to admit students to the public universities. The scores on this test are normally distributed with a mean of 350 and a standard deviation of 60. Mary knows if she wants to be admitted to a certain university, her score should be better than at least 85% of the students who took the test.

a) (6%) Mary takes the test and scores 435. Will she be admitted to that certain university?

b) (7%) If a random sample of 40 students selected, what is the probability that the sample average of test scores is at most 370?

c) (7%) How large a sample size would be required to ensure that, the probability of having the average of 375 for test scores is at least 0.95?
Answer #1

let the scores are random variable x XnNl350 (60)²] we know that it from NCM.02) then. Z = X - M رله ما P(XK 435)=P(x-350 435n = ? P(x<375) - 0.95 P(x-350 375-350 60/ h ) = 0.95 Golsh plek 375-350 -0.95 60/8n corresponding to from 2-table calculator